Find the axis of symmetry...
1. y=-x^2+4x+2
Find the x-intercepts...
1. y=x^2+2x-8
The length of a rectangle is 3cm more than 2 times its width. If the area of the rectangle is 99cm^2, find the dimensions of the rectangle to the nearest thousandth.

Answer by stanbon(75887)

Find the axis of symmetry...
1. y=-x^2+4x+2
Complete the square to get:
-(x^2-4x+4) = y-2-4
-(x-2)^2 = y-6
Axis of symmetry: x=2
-----------------------
Find the x-intercepts...
1. y=x^2+2x-8
Let y=0 and solve for x:
x^2+2x-8=0
(x-2)(x+4)=0
x=2 or x=-4
---------------------
The length of a rectangle is 3cm more than 2 times its width. If the area of the rectangle is 99cm^2, find the dimensions of the rectangle to the nearest thousandth.
Let the width be "x" cm; the length is "2x+3" cm
Area = x(2x+3)
99 = 2x^2+3x
2x^2+3x-99=0
x=[-3+-sqrt(9-4*2*-99)]/4
Positive answer is:
x= 25.31/4
x= 6.325 cm (this is the width)
2x+3 = 15.65 cm (this is the length)
===============
Cheers,
Stan H.